# Responsible AI Question Bank: A Comprehensive Tool for AI Risk Assessment

### v1

@@ 4.2 Human, Societal, Environmental Wellbeing
The RAI Question Bank can be used to assess both positive and negative impacts of the AI system on human, society and the environment. Accordingly, the principle question is defined as "Does the AI system benefit human, society and environment?" (Figure 5). There are a total of 14 sub-questions under the principle question to dive deeper and conduct practical-level risk assessments.
@@
Although the questions pertain to the same risk theme, users may not need to use simultaneously. We have defined appropriate stages for each question, considering the AI development life cycle as mentioned in the previous section. Figure 5 shows example questions and the proposed stages at which users should ask them.
@@ 4.3 Human-centred Values
It accommodates the need for an appropriate level of human oversight and control to prevent AI risks, particularly those associated with human rights. Under the key question, "Does the AI system respect human rights, diversity and autonomy of individuals?", the RAI Question Bank includes 17 sub-questions primarily based on three frameworks: NIST, EU and Microsoft. The questions are categorized into three risk themes: human rights, human oversight and human agency (Figure 6).
@@
Human rights is designed with five dedicated questions to identify risks related to regulatory requirements and ensure that the AI system respects human rights.
Human agency comprises three questions and includes one sub-category, Task allocation. This category assesses how the AI systems and human share tasks, whether the AI system enhances or augments human capabilities, and if there is an appropriate level of control over tasks allocated to AI systems.
Human oversight includes nine sub-questions aimed at assessing AI risks due to the lack or absence of human involvement, with defined processes and mechanisms to govern and control AI. the RAI Question Bank suggests that these questions should be used at the early stage (i.e., planing stage) of the AI lifecycle to establish a solid governance framework early on and support its implementation across the categories.
@@ 4.4 Fairness
The key concept of this principle includes that AI systems should be inclusive and accessible, and should not involve or result in unfair discrimination against individuals, communities or groups. This concept may lead to responsible design and deployment of AI systems minimizing bias and promoting inclusion and fairness (Figure 7). the RAI Question Bank includes 32 questions to thoroughly assess risks associated with this concern.
@@ 4.5 Privacy and Security
Privacy and security concerns can be assessed by 47 questions. Using these questions, categorized into three risk categories—privacy protection, data protection, and data quality management—users can ensure that AI systems are designed with privacy and security in mind from the early stages to protect sensitive and copyrighted data used by the AI systems.
Figure 8 depicts the risk concept map related to privacy and security concerns in the context of AI, along with the key risk question and example subsequent questions.
@@
4.6 Relaiability and Safety
This principle comprises five categories: system performance, system test, system reliability, system resilience, and adverse impact (Figure 9). Together, these categories encompass a total of 42 risk questions designed to thoroughly assess and ensure the reliability and safety of AI systems.
@@ 4.7 Transparency and Explainability
The RAI Question Bank includes 32 questions for three risk categories (Figure 10): Transparency, Explainability, and Communication. This principle specifically underscores the responsibility of the AI model and system providers towards the users to ensure that AI systems are transparent, understandable, and effectively communicated.
@@
Communication encompasses questions that evaluate the effectiveness of the information exchange between the AI system and stakeholders including users. This category ensures that the system provides timely, accurate, and useful information to them, facilitating informed decision-making and fostering a positive user experience.
@@ 4.8 Contestability
Contestability refers to the ability to challenge and question AI decisions, ensuring users and affected parties can seek redress or have decisions reviewed. Most existing RAI frameworks do not explicitly address or include this principle. We have observed that the AIA and NIST frameworks include a few questions related to this principle. Other frameworks tend to address this concept implicitly under their Transparency or Fairness principles.
We identified four questions to support this principle, focusing on human interfaces and the right to appeal. These questions address mechanisms for users to contest AI decisions, the processes in place for reviewing contested decisions, and the transparency of these processes. By incorporating contestability measures, AI systems promote accountability and provide users with fair means to address potential biases or errors in AI outputs.
@@ 4.9 Accountability
To address risk associated with this principle, the RAI Question Bank includes 57 questions distributed into five categories: Auditability, Trade-offs analysis, Redressibility, Accountability framework, and AI management (Figure 12). This principle encompasses a more extensive set of questions compared to other principles due to recent additions aligned with emerging AI policies such as the EU AI Act and ISO AI management. The heightened focus on comprehensive AI quality and management within these regulations has generated such additional accountability-related practices and risk questions.
@@
Auditability ensures that AI systems are designed with mechanisms that allow for comprehensive audits, enabling stakeholders to trace and verify decisions made by the AI. This category includes 11 questions that assess the ability to monitor and review AI operations and outcomes, ensuring transparency and trustworthiness.
Trade-offs analysis focuses on the decision-making processes within AI systems, particularly how they balance competing objectives and values. This includes questions that identify relevant interests and values implicated by the AI system, potential trade-offs between them, and the consideration of ethical implications and potential conflicts between different goals. This ensures that AI systems make balanced and fair decisions while also aiming to understand potential risks associated with the lack or absence of management of trade-offs.
Redressibility addresses the mechanisms in place for correcting or mitigating harm caused by AI systems. It includes two questions that assess the existence and effectiveness of processes for users to seek redress in case of any harm or adverse impact and provide information to the users (or third parties) about opportunities for redress.
@@ 5.1 RAI risk assessment in Scientific Research
Case study overview. We have evaluated the ethical risks associated with eight AI projects, designated PR1 to PR8, within a national research organization in Australia. These projects primarily focus on scientific research and incorporate AI as a component of broader research initiatives. Each project has distinct goals (see Table 2) and operates across various sectors such as food security and quality, health and well-being, and sustainable energy. Given that these projects are in their early stages, they provide a valuable opportunity to integrate best practices in responsible AI for future development.
Project code Project objective/context
PR1 Using AI to design & optimisation of real-world machinery and processes.
PR2 Understanding the casual relationships between climate systems and social phenomenon such as pandemics, conflicts etc. using AI-assisted technologies.
PR3 Using ML/AI for drug discovery.
PR4 Using ML/AI for bio security applications.
PR5 Using AI-assisted tools to identify pulse varieties for food replacement.
PR6 Using AI for coastal forecasting based on water quality models.
PR7 Developing AI based tools and technologies for managing high penetration of Distributed Energy Resources (DER) in the electricity network.
PR8 Developing suite of tools including AI to trace final meat cuts back to the individual animal.
Table 2: Assessed projects and context.
The hierarchy of the project teams was relatively flat, consisting of a Project Lead who manages the entire research project, multiple Work Package Leads who oversee sections of the research project, and Researchers who report to the Work Package Leads.
We conducted this case study to primarily serve three purposes concerning the RAI Question Bank:
  1. 1.
    To test the feasibility of the RAI Question Bank in assessing ethical AI risks, which is one of the key use cases of the RAI Question Bank.
  2. 2.
    To identify the challenges of using the question bank when applied in a real-world project.
  3. 3.
    To test the question bank in a specific context—in this case, scientific research using AI.
@@
The role of RAI–QB. Prior to conducting the risk assessment, we developed a dedicated risk register template that includes essential risk assessment components such as risk ID, risk category, risk description, risk causes, existing mitigation measures, risk owners, and interview questions. the RAI Question Bank was integral to this template, serving as the foundation for developing risk assessment categories, different assessment levels, and interview questions. For instance, the risk assessment register adhered to the risk concept map for RAI, which includes 26 risk themes (see Figure  3). Additionally, it utilized the three question levels and corresponding questions from the RAI Question Bank to formulate interview questions for each of these risk themes.
@@
Risk assessment. The first round of risk assessment sessions was scheduled between January and February 2023. Each session involved an interview lasting approximately 1.5 hours. The participants included a risk assessor, 1-2 observers, and 1-4 interviewees. Project leads of all 8 projects participated in the interviews, and in most cases, the work package leads also participated. In a few instances, researchers participated as interviewees as well.
@@
In the first round, we focused on the high-level (level 1 and 2) questions of the RAI Question Bank for each AI principle. The interviews were recorded and transcribed using a paid service, and the transcriptions were rechecked for accuracy. After an initial data analysis, another series of interviews was conducted (round 2) a few months later. The main motivation behind the second round of interviews was to discuss more in-depth information on the risks identified, primarily using the level 3 questions of the RAI Question Bank.
The collective interview data were analyzed to identify and assess the risks associated with AI ethics in these 8 projects. For risk level calculations, we used a well-known 3x3 risk assessment matrix and an impact and probability equation [25].
Discussion on effectiveness of the RAI Question Bank in identifying risks. The question bank’s structured approach, combining high-level principles with the different-level questions, has proven to be effective in uncovering potential risks across diverse AI projects. By mapping questions to different development phases and incorporating feedback from users, we have ensured that the question bank addresses a broad range of risk factors and provides actionable insights.
@@
Through the first round of interviews, we identified Accountability emerges as the principle with the highest combined medium to high risk. This is mainly to do with missing a proper responsible distribution documentation and lack of traceability within projects. Transparency and explainability also pose significant risks, primarily due to insufficient clarity in explaining AI outcomes. While accuracy is crucial in scientific discovery, it can result in less understandable results. Implementing additional measures is advised to ensure that stakeholders have a clear understanding of AI outcomes, which is vital for informed decision-making. Conversely, Human-centred values, Human, societal and environmental wellbeing and Contestability are recognized as relatively low-risk principles. These areas have demonstrated better alignment with ethical practices and project requirements, suggesting fewer immediate concerns.
The key findings and risks for each project were further discussed in the second round of interviews, confirming that the RAI Question Bank effectively highlights critical risk areas and provides a solid foundation for risk management.
Discussion on feasibility to use the RAI Question Bank for assessing ethical AI risks. The feasibility of using the RAI Question Bank for assessing ethical AI risks was demonstrated through its structured approach to risk assessment. The Question Bank provided a comprehensive framework that facilitated navigation through various risk areas, ensuring both breadth (covering the 8 AI principles) and depth (addressing different levels of questions) in the assessment process.
The structured questions prompted interviewees to consider different aspects of their projects that they had not previously contemplated. For instance, one of the Project Leads in PR3, which focuses on AI for drug discovery, had not considered the risks associated with third-party university agreements on intellectual property ownership, particularly regarding sharing chemical formulas stored in university repositories, until prompted by the RAI Question Bank questions. The structured and comprehensive nature of the RAI Question Bank helped uncover a wide range of ethical considerations, enhancing the overall risk assessment process.
@@
Discussion on practical challenges of using the RAI Question Bank in real world projects. Durging this case study, several practical challenges were encountered when using the RAI Question Bank to assess ethical AI risk in real-world projects. Firstly, the stage of the project significantly influences the risk assessment process. For example, projects at the inception phase may have a limited understanding of model accuracy, which impacts the evaluation of the principle of reliability and safety under the sub category - system performance (see Figure 9).
Secondly, the effectiveness of the interviews depends heavily on the interviewees. In our context, the flat hierarchy meant that project leads and work package leads, who were also researchers, had comprehensive knowledge of their projects. Therefore, they were able to answer all the question from level 1 to 3. However, we acknowledge that this may not be the case in many commercial or open-source collaborative projects, where individuals may only have knowledge specific to their roles.
Time allocation also posed a challenge. The initial 1.5-hour sessions were insufficient, necessitating a second round that focused on the ethical principles found comparatively more important to the project. This iterative approach to risk assessment proved more effective, suggesting that less critical principles could be de-prioritized. Additionally, equal time allocation for each principle did not work well. Participants spoke more extensively about principles they were familiar with, such as privacy, especially when their projects were directly related to those principles.
To address these challenges when using the RAI Question Bank in ethical AI risk assessment, it would be beneficial to prioritize ethical principles based on publicly available information or through a pre-screening process, ensuring a more focused and efficient assessment.
Discussion on using the RAI Question Bank in the context of AI related scientific research projects. Using the RAI Question Bank in the context of AI-related scientific research projects presents several unique challenges. Many of these projects are not yet at the stage where they engage with real-world users, making it difficult to assess risk levels based on the ethical AI principles’ definitions. This disconnect means that some of the risk definitions and questions may not be directly applicable to the specific context of these scientific research projects, until they reach the point of commercialisation.
The context of scientific research often differs from common AI applications, which can sometimes lead to gaps between the RAI Question Bank principles and the needs of these projects. While the RAI Question Bank offers a thorough framework for assessing ethical risks, its questions are generally aimed at common real-world issues. Scientific research, however, focuses on solving complex scientific problems, which might not always fit neatly with the RAI Question Bank’s framework. Despite this, the RAI Question Bank’s flexibility allows for adjustments and can be tailored to better fit the specific needs of scientific research projects. Moreover, this proves, the RAI Question Bank can be used in various different risk assessment contexts.
@@
Feedback for improvement. Following the risk assessment process, stakeholders provided valuable feedback. One key point was that the number of questions in the RAI Question Bank sometimes seemed excessive relative to the level of completeness needed for the assessment. While the entire question bank may useful for ethical AI assessments in mission-critical or life-critical systems, a subset of questions could be sufficient for systems with lower risk levels.
@@
Another point of feedback was the relevance of the questions. Although the RAI Question Bank questions were generally useful for identifying risks, not every question was applicable at every stage of the project. This was recognized as a manageable challenge but highlighted the need for improvement.
Based on this feedback, stakeholders recommended enhancing the RAI Question Bank structure by incorporating a dimension that accounts for the development stage of the project. Tailoring questions to the specific stage of each project would enable more accurate and contextually relevant risk identification. This enhancement is expected to improve the applicability, efficiency, and effectiveness of the RAI Question Bank, contributing to more robust and comprehensive AI risk management practices.

### v2

@@ 4.2 Human, Societal, Environmental Wellbeing
The rapid diffusion of AI has touched nearly every aspect of our lives. Many AI applications are developed using human-contributed data, directly interact with users (e.g., chatbots), or perform tasks on behalf of humans. However, the rise of AI is accompanied by concerns such as job displacement and harmful outcomes, both for individuals and society at large [43, 44, 45]. Specifically, the widespread adoption of generative AI can introduce risks of toxicity, such as promoting hate, violence, or causing offense [46]. While the societal harms of AI are sometimes overlooked, its environmental impacts such as carbon emissions, energy consumption, and sustainable ecosystems have been widely discussed across various industries [47]. These concerns underscore the critical challenge of designing AI systems that support human health and well-being [48].
The RAI Question Bank can be used to assess both positive and negative impacts of the AI system on human, society and the environment. The key question is defined as "Does the AI system benefit human, society and environment?" (Figure 4). There are a total of 14 sub-questions under the principle question to dive deeper and conduct practical-level risk assessments.
@@
Although the questions pertain to the same category, users may not need to use simultaneously. We have defined appropriate stages for each question, considering the AI development life cycle as mentioned in the previous section. Figure 4 shows example questions and the proposed stages at which users should ask them.
@@ 4.3 Human-centred Values
With the increasing reliance on AI-based decision-making systems in recent years, there is a growing demand for ethical AI, which is often perceived as requiring human oversight of automation [49]. Although AI can significantly enhance performance compared to manual decision-making processes, over-reliance on these systems can result in negative consequences, particularly in complex or high-stakes situations where human oversight is essential [50]. Therefore, emphasizing human agency and oversight has become a key requirement for the effective implementation of AI applications [51, 52].
To cope with these concerns, RAI should consider an appropriate level of human oversight and control to prevent AI risks, particularly those associated with human rights. Under the key question, "Does the AI system respect human rights, diversity and autonomy of individuals?", the RAI Question Bank includes 17 sub-questions primarily based on three frameworks: NIST, EU and Microsoft. The questions are categorized into three: human rights, human oversight and human agency (Figure 5).
@@
Human rights is designed with five dedicated questions to identify risks related to regulatory requirements and ensure that the AI system respects human rights. This category emphasizes the impact of technology design such as AI system on human autonomy [53]. Human autonomy refers to the ability to willingly choose to be either dependent or independent [54]. Ensuring that AI systems uphold human autonomy is crucial for fostering trust and preventing unintended manipulation or coercion.
To maintain effective human oversight and intervention, there should be a clear delineation of responsibilities between human operators and AI systems [55]. Human agency category comprises three questions and includes one sub-category, Task allocation. This assesses how the AI systems and human share tasks, whether the AI system enhances or augments human capabilities, and if there is an appropriate level of control over tasks allocated to AI systems.
Human oversight includes nine sub-questions aimed at assessing AI risks stemming from insufficient human involvement, with processes and mechanisms in place to govern and control AI. Additionally, it recommends direct and immediate solutions to mitigate negative impacts, such as a stop button or procedure to safely abort operations when necessary, as commonly required by AI regulations such as the EU AI Act [56]. We suggest that these questions should be used at the early stage (i.e., planing stage) of the AI lifecycle to establish a solid governance framework early on and support its implementation across the categories.
@@ 4.4 Fairness
Fairness has attracted significant attention from researchers seeking to address the negative consequences of AI, such as unfair biases that can lead to discrimination or poor system performance [57]. AI systems should be inclusive, accessible, and free from unfair discrimination against individuals, communities, or groups [58]. Adopting such an approach promotes responsible AI design and deployment, minimizing bias and fostering inclusion and fairness. Bias can emerge at various stages of AI systems, including in the data, models, or outputs. As AI systems learn patterns from input data to generate results, it is essential to address fairness throughout the entire AI life cycle [59]. This requires ongoing efforts to identify, monitor, and control bias at every stage.
The RAI Question Bank includes 32 questions to thoroughly assess risks associated with this concern (Figure 6).
@@
There are various techniques and metrics to support AI fairness. Disparate impact, equal opportunity, and statistical parity are commonly used to assess whether a system was designed without discriminatory intent, ensuring equal opportunities and statistically similar outcomes for all stakeholder groups [60, 61, 62]. However, achieving fairness often involves trade-offs, as optimizing for one fairness metric can sometimes lead to reduced performance in another area or conflict with other ethical principles [63].
@@ 4.5 Privacy and Security
AI systems rely heavily on human-contributed data, raising significant concerns about personal information, sensitive data, privacy violations, and copyright issues [64]. Several privacy-preserving solutions, such as federated learning [65] and differential privacy [66], have been proposed, though no perfect solution exists. This highlights the importance of implementing privacy and security measures during the design phase—or earlier—and maintaining them throughout the entire life cycle to minimize technical debt [67, 68].
AI systems must ensure personal data privacy across their life cycle, including data traceability and protection against unauthorized access [69]. Figure 7 illustrates the RAI concept map related to these concerns, along with key risk questions and example follow-up questions. Privacy and security concerns can be assessed through 47 questions, categorized into three risk areas: privacy protection, data protection, and data quality management. These questions help ensure that AI systems are designed with privacy and security in mind from the early stages, safeguarding sensitive and copyrighted data throughout their operation.
@@
4.6 Reliability and Safety
Reliability and safety are key components of a system’s trustworthiness [70]. AI systems should be designed, developed, and tested with a focus on human safety, particularly in safety-critical domains such as healthcare, where system failures can directly cause patient harm [71]. Reliability and safety go beyond achieving high accuracy. According to the EU AI Act, high-risk AI systems must ensure comprehensive robustness, including appropriate accuracy metrics, resilience to errors, faults, or inconsistencies, risk management, and fail-safe mechanisms. These systems should also be tested against predefined metrics and probabilistic thresholds aligned with their intended purpose [21]. A holistic approach is essential to meeting these requirements.
We have identified five categories for these concerns: system performance, system test, system reliability, system resilience, and adverse impact (Figure 8). Together, these categories encompass a total of 42 risk questions designed to thoroughly assess and ensure the reliability and safety of AI systems.
@@ 4.7 Transparency and Explainability
There is broad consensus that AI systems, particularly deep learning models, are often opaque and difficult for humans to understand—commonly referred to as the black box problem [72]. This lack of transparency has contributed to diminished user trust, driving increased demand for transparent and explainable AI in the RAI space [73]. As a result, most AI governance frameworks and regulations now include Transparency and Explainability as essential principles for developing and deploying trustworthy AI systems [21, 18, 74].
Traditionally, AI transparency focused on algorithms, aiming to demonstrate how decisions were made [75]. However, recent frameworks call for more comprehensive information about AI systems, including specifications such as the intended purpose, data sources, system performance, required training resources (e.g., computing power and time), and limitations (e.g., unmitigated risks) [21]. Additionally, interpretability should be prioritized as a key design element to ensure impartiality, robustness, and meaningful insights from system variables [73].
To address these concerns, the RAI Question Bank includes 32 questions for three risk categories (Figure 9): Transparency, Explainability, and Communication. This principle specifically underscores the responsibility of the AI model and system providers towards the users to ensure that AI systems are transparent, understandable, and effectively communicated.
@@
Communication encompasses questions that evaluate the effectiveness of the information exchange between the AI system and stakeholders including users. Providing information to non-technical audiences may require managing ambiguities, but the information should be accessible for society, policy makers and any other stakeholders. [73]. This category ensures that the system provides timely, accurate, and useful information to them, facilitating informed decision-making and fostering a positive user experience.
@@ 4.8 Contestability
Contestability refers to the ability to challenge and question AI decisions, ensuring users and affected parties can seek redress or have decisions reviewed [76]. Most existing RAI frameworks do not explicitly address or include this principle. We have observed that the Canada AIA and Australia NSW frameworks include a few questions related to this principle. Other frameworks tend to address this implicitly under their Transparency or Fairness principles. This issue has resulted in a lack of guidance on designing AI systems that support contestation.
In human decision-making, it is generally easy to obtain the reasoning behind a decision from the decision-maker through internal or external reviews and complaints mechanisms. However, AI-based decision-making is inherently complex and often operates at scale, making it difficult for stakeholders to access the decision process and raise objections [77]. It underscores that AI contestability should be incorporated early in the development process, with mechanisms embedded within the AI system through an interface. This Contestability by Design concept can also help AI systems achieve fairness by making them contestable for all stakeholders [78].
In line with these considerations, we identified four questions to support this principle, focusing on human interfaces and the right to appeal. These questions address mechanisms for users to contest AI decisions, the processes in place for reviewing contested decisions, and the transparency of these processes. By incorporating contestability measures, AI systems promote accountability and provide users with fair means to address potential biases or errors in AI outputs.
@@ 4.9 Accountability
Accountability in the RAI space is defined as the principle that AI actors should be held responsible for the proper functioning of AI systems [74]. It extends beyond merely defining roles and responsibilities of AI actors to include associated tasks and activities, such as ensuring traceability (across datasets, processes, and decisions) and implementing essential management frameworks, including quality, risk, leadership, and competency management [79, 80, 81].
Accountability challenges include unclear responsibility for AI outcomes and the absence of mechanisms to address unintended consequences [82]. Accountability is closely tied to other RAI principles, such as Fairness, Privacy, and Transparency, as it provides fundamental management frameworks that support their implementation. As it plays a pivotal role, the lack or absence of an appropriate level of accountability in RAI can critically impact trust in AI systems and lead to misuse, unethical practices, or harmful outcomes.
To address these risks, the RAI Question Bank includes 57 questions distributed into five categories: Auditability, Trade-offs analysis, Redressibility, Accountability framework, and AI management (Figure 11). This principle encompasses a more extensive set of questions compared to other principles due to recent additions aligned with emerging AI policies such as the EU AI Act and ISO AI management. The heightened focus on comprehensive AI quality and management within these regulations has generated such additional accountability-related practices and risk questions.
@@
Auditability ensures that AI systems are designed with mechanisms that allow for comprehensive audits, enabling stakeholders to trace and verify decisions made by the AI [83]. This category includes 11 questions that assess the ability to monitor and review AI operations and outcomes, ensuring transparency and trustworthiness.
Trade-offs analysis focuses on the decision-making processes within AI systems, particularly how they balance competing objectives and values [84]. This includes questions that identify relevant interests and values implicated by the AI system, potential trade-offs between them, and the consideration of ethical implications and potential conflicts between different goals. This ensures that AI systems make balanced and fair decisions while also aiming to understand potential risks associated with the lack or absence of management of trade-offs. An example of trade-off analysis is an accuracy-energy experiment, which demonstrates that in some cases, a small reduction in accuracy can result in significant energy savings [85]. Another important trade-off for AI systems is between data quality and data availability.
Redressibility addresses the mechanisms in place for correcting or mitigating harm caused by AI systems [86]. It includes two questions that assess the existence and effectiveness of processes for users to seek redress in case of any harm or adverse impact and provide information to the users (or third parties) about opportunities for redress.
@@ 5.1 RAI Risk Assessment in Scientific Research
Case study overview. We have evaluated the ethical risks associated with eight AI projects, designated PR1 to PR8, within a national research organization in Australia. These projects primarily focus on scientific research and incorporate AI as a component of broader research initiatives. Each project has distinct goals and operates across various sectors such as food security and quality, health and well-being, and sustainable energy. Given that these projects are in their early stages, they provide a valuable opportunity to integrate best practices in responsible AI for future development. The hierarchy of the project teams was relatively flat, consisting of a Project Lead who manages the entire research project, multiple Work Package Leads who oversee sections of the research project, and Researchers who report to the Work Package Leads.
@@
Preparation— Development of a risk register template. Prior to conducting the risk assessment, we developed a dedicated risk register template that includes essential risk assessment components such as risk ID, risk category, risk description, risk causes, existing mitigation measures, risk owners, and interview questions. The RAI Question Bank was integral to this template, serving as the foundation for developing risk assessment categories, different assessment levels, and interview questions. The risk assessment register adhered to the risk categorization which includes 26 themes (see Figure  3). Additionally, it utilized the three question levels and corresponding questions from the RAI Question Bank to formulate interview questions for each of these themes.
@@
Risk assessment— Engagement with project teams. The first round of risk assessment sessions was scheduled between January and February 2023. Each session included a 1.5-hour interview, with 10 minutes allocated for introduction and wrap-up. The participants included a risk assessor, 1-2 observers, and 1-4 interviewees. Project leads of all eight projects participated in the interviews, and in most cases, the work package leads also participated. In a few instances, researchers participated as interviewees as well.
@@
In the first round, we focused on the high-level (level 1 and 2) questions of the RAI Question Bank for each AI principle. The interviews were recorded and transcribed using a paid service, and the transcriptions were rechecked for accuracy. Following the initial data analysis, a second round of interviews was conducted a few months later. It aimed to explore the identified risks in greater depth, focusing primarily on the level 3 questions from the RAI Question Bank.
@@
Results— Effectiveness and feasibility. The question bank’s structured approach, combining high-level principles with the different-level questions, has proven to be effective in uncovering potential risks across diverse AI projects. The structured questions prompted interviewees to consider different aspects of their projects that they had not previously contemplated. For instance, one of the Project Leads in PR3, which focuses on AI for drug discovery, had not considered the risks associated with third-party university agreements on intellectual property ownership, particularly regarding sharing chemical formulas stored in university repositories, until prompted by the RAI Question Bank questions. Through the first round of interviews, we identified Accountability emerges as the principle with the highest combined medium to high risk. This is mainly to do with missing a proper responsible distribution documentation and lack of traceability within projects. Transparency and explainability also pose significant risks, primarily due to insufficient clarity in explaining AI outcomes. Conversely, Human-centred values, Human, societal and environmental wellbeing and Contestability are recognized as relatively low-risk principles. These areas have demonstrated better alignment with ethical practices and project requirements, suggesting fewer immediate concerns. The key findings and risks for each project were further discussed in the second round of interviews, confirming that the RAI Question Bank effectively highlights critical risk areas and provides a solid foundation for risk management.
The feasibility of using the RAI Question Bank was demonstrated through its structured approach. It provided a comprehensive framework that facilitated the exploration of diverse risk areas, ensuring both breadth (covering the eight AI principles) and depth (addressing multiple question levels). This structure enabled the efficient creation of a risk register template and formulation of interview questions. We analyzed project contexts and selected 3 to 5 questions per AI principle to guide the interviews. The risk assessments incorporated the risk register template, a standard 3x3 risk matrix, and calculations for impact and probability to determine risk levels. By applying the question bank across eight projects, we confirmed its value as a focal resource, showing that it integrates smoothly into traditional risk workflows without requiring significant adjustments.
@@
Discussion— Practical challenges in real-world contexts. Durging this case study, several practical challenges were encountered when using the RAI Question Bank to assess ethical AI risk in real-world projects. Firstly, the stage of the project significantly influences the risk assessment process. For example, projects at the inception phase may have a limited understanding of model accuracy, which impacts the evaluation of the principle of reliability and safety under the sub category - system performance (see Figure 8). Secondly, the effectiveness of the interviews depends heavily on the interviewees. In our context, the flat hierarchy meant that project leads and work package leads, who were also researchers, had comprehensive knowledge of their projects. Therefore, they were able to answer all the question from level 1 to 3. However, we acknowledge that this may not be the case in many commercial or open-source collaborative projects, where individuals may only have knowledge specific to their roles.
@@
Feedback and improvement. Following the risk assessment process, stakeholders provided valuable feedback. One key point was that the number of questions in the RAI Question Bank sometimes seemed excessive relative to the level of completeness needed for the assessment. While the entire question bank may useful for ethical AI assessments in mission-critical or life-critical systems, a subset of questions could be sufficient for systems with lower risk levels.
@@
Another point of feedback was the relevance of the questions. Although the risk questions were generally useful for identifying risks, not every question was applicable at every stage of the project. This was recognized as a manageable challenge but highlighted the need for improvement.
Based on this feedback, we reviewed and refined the question bank, mapping questions to different development phases. This enhancement enables a more efficient and effective risk identification process, aligning questions with each project’s specific stage. By tailoring questions to the development stage, potential risks can be identified more accurately and in a manner more applicable to the project’s context. This not only enhances the relevance and usefulness of the question bank but also optimizes resource allocation and efforts in assessing AI risks. This refinement addresses the challenge of assessing risk levels in early-stage projects with limited model accuracy insights and minimal real-world user engagement. By aligning questions with specific development stages, we ensure that the questions are more applicable and actionable throughout the project life-cycle. Additionally, this approach mitigates the impact of time constraints by prioritizing the most relevant ethical principles based on the project’s current phase, ultimately leading to a more efficient and focused risk assessment process.